### v2

# Mathematical Politics

Abstract
Politics today is largely about the art of messaging to influence the public, but the mathematical theory of messaging — information and communication theory — can turn this art into a precise analysis, both qualitative and quantitative, that enables us to gain retrospective understandings of past political events and to make forward-looking future predictions.
Introduction
Over the years I have observed that some of my colleagues in mathematical science community are oblivious to what is happening in politics. Perhaps this is a healthy attitude, given all that is taking place around the world today, but for others who are more in tune with politics, it might be a revelation that the power of mathematical analysis stretches into politics to make concrete statements. Mathematical politics, or ‘political mathematics’, is not about politics taking place within the mathematical science community (about which those colleagues do have much to say!). Rather it concerns mathematical modelling that enables us to understand political events, such as elections or referendums, and in some cases allows us to make future predictions. [^1] I am not concerned here with mathematical approaches to social choice in political economy (Arrow 1963, Taylor 1971), nor with any other areas of mathematical economics that may have implications in policy choice. The term ‘mathematical politics’ has indeed been used once in the context of Arrow’s impossibility theorem (Samuelson 1967). Today’s politics, however, has diverged away from the traditional idea that mathematical economics might assist politicians in designing better economic policies. Rather, with the advances in communication technology, today’s politics has reduced largely to the art of messaging, often lacking in substance for ideas to tackle real societal challenges, and this is where the science of information, communication, and control can play significant roles, both in positive and negative ways.
The word politics carries multiple meanings, but broadly speaking I have in mind here the activities of the government or lawmakers to try to influence the way a country is governed. A country can be replaced by an organisation, in which case we will be speaking of an organisational politics, but let me stay with the governance of a country in a broadly democratic system. Then to implement a policy, one must either convince the government to take your advice, or else convince the voters to allow you form a government. From this point of view, politics boils down to convincing others through messaging. If we want a scientific approach to this process, then we must understand how people digest messages and extract information contained therein. This is where we can apply powerful mathematical techniques of information and communication theory with the hope of making a difference, and that is the topic I would like to outline in what follows. I begin by discussing closely-related ideas in financial modelling, which provides a context in which it is easier to understand the role of information in society. Then I proceed to discuss the calculus for elections, and explore other ideas that are important in political messaging.
A question arising from financial modelling
Norbert Wiener, arguably one of the greatest mathematicians to have been born in America, made the point in the late 1940s that society can only be understood through the lens of communication theory (Wiener 1950), because communication is the basis of society. Although Wiener made a broad range of predictions on the future forms of society, many of which, including one on the environmental challenge (box below), became reality, he did not implement any concrete mathematical model to study any particular social system.
From The Human Use of Human Beings (Wiener 1950)
“This is partly the result … of an increased mastery over nature which, on a limited planet like the earth, may prove in the long run to be an increased slavery to nature. For the more we get out of the world the less we leave, and in the long run we shall have to pay our debts at a time that may be very inconvenient for our own survival. … We have modified our environment so radically that we must now modify ourselves in order to exist in this new environment.”
Neither did anyone else implement Wiener’s vision in the fifty years after his book of 1950 in the context of social science, although the mathematical concepts underpinning information and communication theory developed rapidly in areas such as coding theory, signal processing, and stochastic filtering, all of which have significant practical applications in various areas of engineering and technology.
Only in the past two decades, a new and unexpected application emerged that embodies Wiener’s philosophy in the context of financial modelling. In the financial industry, the pricing and risk management of financial contracts was a big issue twenty to thirty years ago. These financial contracts, known as contingent claims, entail random payouts that are dependent on the future prices of one or more underlying assets. An example of such an asset is a stock. The pricing issue is then not concerned with the price of the stock, which is determined by the market, but the pricing of a contract whose payout is a function of the future stock price.
Following the tradition of mathematical modelling, one thus requires a model for the price of the stock, from which the prices of contingent claims on that stock can be worked out. This is natural inasmuch as we observe the daily ups and downs of the stock price, so we construct a model for what we see, in the form of a solution to a stochastic differential equation. But is it really the right approach, if we want to obtain logically motivated asset-price models?
Understanding information flow is the key
The issue here is that the daily movements of asset prices result as a consequence of the daily revelation of information. Yet, if we start out from a model for the asset price, it is not possible to connect that model to the information that generated the price movements in the first place. In other words, the causal relation between the flow of information and the price dynamics is lost. It is then inevitable that the chosen model for the stock price will have little to do with the actual price movements. So there has to be a big leap of imagination to do better. But how?
For a long time, economists have understood that the main driving factor of price movements in financial markets is information. However, unlike price movements of a stock that can be measured as a time series, information, such as news bulletins, does not come in the form of a numerical time series. The intuition about the impact of information on the price is clear enough — we read a news article about global protests of Tesla cars and we expect the Tesla stock price to drop. But this news article is presented in the form of words and sentences. So how can we model such information mathematically?
The answer is to rely on the notion of a ‘random variable’ in probability. A random variable assigns a number to an outcome of chance, and is an indispensable tool for modelling the statistics of random phenomenon. An example is the outcome of coin toss, which is either a head or a tail. If we wish to apply statistical analysis to the outcomes of a fair coin, such as the mean of the outcome, we encounter an issue because we cannot average a head and a tail — there is no ‘Cecrops’ face of a coin. But if we assign the numerical value zero to a tail and one to a head, say, which is what a random variable does, then we can meaningfully say that the average outcome of a fair coin is a half.
The idea is to extend this notion to general information that we digest. For this purpose, the first step is to identify the relevant quantity of interest contained in a message, and associate numerical values to label different realisations of this quantity. For example, in a financial investment, the question an investor would ask is: What is the return? Needless to say, no one is actively transmitting this information, unlike in a conventional communication channel. Nevertheless, we can interpret the quantity of interest as the signal component of a message in communication theory. Of course, information is rarely perfect in financial markets — no one knows the future price of an asset. The imperfection of information, representing rumours, speculations, or misinterpretations, can then be modelled in the form of noise in a communication channel. Putting these ideas into a simple formula, the acquisition of noisy information on the future payout $X$ of an investment can be translated into the detection of the value of the noisy information
$\xi=\sigma X+\epsilon\,$,
where $\epsilon$ represents noise, modelled by a random variable independent of $X$, and I have also inserted the signal-to-noise ratio $\sigma$.
In the case of a financial investment, the modelling of the signal is straightforward because the quantity of interest, namely the return, is already expressed numerically. In general this is not the case, and this is where one has to be creative in assigning numerical values to different realisations of the quantity of interest as a random variable. For example, if one is concerned with making a choice among $N$ alternatives, then these alternatives can be labelled by $N$ numbers $x_{1}$, $x_{2}$, $\cdots$, $x_{N}$. Let $p_{k}$ be the current likelihood of alternative $x_{k}$ being selected. Then we have a model for the signal $X$. For a binary choice we thus have $X=x_{1}$ or $X=x_{2}$. Any speculative information (noise) that suggests that the correct choice is $X=x_{2}$ is represented by the random variable $\epsilon$ taking a large value — the larger the value is, the stronger the suggestion is, and conversely smaller values of $\epsilon$ represents speculations towards the alternative $X=x_{1}$. If there are many such speculative or misguided opinions, then the law of large number suggests that $\epsilon$ may be modelled by a zero-mean Gaussian random variable. An information model is then constructed.
Although the mathematics here is simple, the idea it conveys is far reaching, because it goes against our instincts for mathematical modelling: Rather than modelling the price dynamics that we do observe in numbers, we model news items, company reports, central bank announcements, and so on, that we might read on the Internet. So does it work? In financial modelling the proof of the pudding is in the ‘calibrating’ – the process of adjusting model parameters to fit the market prices of contingent claims. For a practical application in financial modelling, though, the arrival of information must be modelled using a time series, for example,
$\xi_{t}=\sigma Xt+\epsilon_{t}\,$,
where we let the noise process $\{\epsilon_{t}\}$ be modelled by a Brownian motion. When an information model of this kind is applied to generate models to price associated contingent claims, the fit to market data was better than the ones used by many of the mainstream investment banks, importantly with fewer model parameters (Mackie 2011). Of course, it is not possible to accurately model the complicated way in which information is spread and digested in the market, but by modelling the root of the price dynamics, one can generate rather realistic models for price processes. As George Box famously put it, ‘all models are wrong’; but these information models are certainly useful. We called this framework an information-based asset pricing (Brody et al. 2008).
Innovation: the arrival of new information
When we digest a ‘news’ article on the Internet or in the papers, often much of the contents in the article is not new. News editors often modify an older article to include the latest developments so that the article is reasonably self-contained. This is indeed quite common — when we acquire information, typically it contains both the known and the unknown. Importantly, though, we tend to respond to the unknown bits because they represent the arrival of genuinely new information.
In communication theory there is a standard procedure to isolate the arrival of new information from an information-providing time series $\{\xi_{t}\}$. To understand this, consider the information gathered over the time interval $t,t+\mbox{$\rm d$}t]$, which is represented by the increment $\mbox{$\rm d$}\xi_{t}$. If we write ${\mathbb{E}}_{t}[-]$ for the expectation conditional on the information available up to time $t$, then the predicted arrival of new information at time $t$ is given by ${\mathbb{E}}_{t}[\mbox{$\rm d$}\xi_{t}]$. It then follows that the discrepancy between the predicted and realised increments
$\mbox{$\rm d$}W_{t}=\mbox{$\rm d$}\xi_{t}-{\mathbb{E}}_{t}[\mbox{$\rm d$}\xi_{t}]$
represents the surprise element, that is, the arrival of new information. The process $\{W_{t}\}$ defined by this relation, which in itself is also a Brownian motion but is rather distinct from the noise, is known as the ‘innovation process’. According to Thomas Kailath, both the concept and terminology were introduced by Wiener (Kailath 1968).
The significance of the innovation process in financial modelling and beyond is the following. The fundamental theorem of asset pricing asserts that the price at time $t$ of an asset entailing a future payout $X$ is given by the expectation ${\mathbb{E}}_{t}[X]$, suitably discounted, in a certain probability measure known as the ‘risk-neutral probability’, conditional on the information available up to time $t$. In the conventional approach to mathematical modelling, as indicated earlier, one specifies a model for the asset price in the form of a stochastic differential equation, driven by a Brownian motion that represents market uncertainties, and assumes that the solution to this equation generates market information, as well as the future return $X$. The hidden implication, not stated in textbooks, is that markets are assumed to be driven by uncertainties.
Markets have views on the future return of an asset, and they are not based on which stochastic differential equation a modeller might have used. These views change in response to new information, leading to dynamics; whereas in the traditional approach, the assumption is that markets obtain information through observing the asset price dynamics. The causality structure is therefore flipped around. This is why in the information-based approach we model the return $X$ along with how information flows to inform markets about the possible values of this return. The price process is then generated through an expectation ${\mathbb{E}}_{t}[X]$, not by solving a differential equation. But having deduced a model for the asset price in this way, one can then ask which differential equation it satisfies. The result reveals that the price process is driven by the innovation process $\{W_{t}\}$, that is, the arrival of new information, and not by market uncertainties $\{\epsilon_{t}\}$; and this is the significant fact about the information-based approach that extends beyond mere financial modelling.
Structural modelling of electoral competition
There are many systems whose dynamics are driven by information, not merely in financial markets. With the advances in artificial intelligence and increased concerns on the impacts of disinformation in society, it will be irresponsible for scientific communities to sit back and pretend that these issues do not concern them, especially when they are equipped with the powerful machinery of the calculus of information. Since the primary concern of the impact of disinformation lies in politics and democratic processes, we need a modelling framework for these processes. The information-based asset pricing theory tells us exactly how this can be done.
As an example, consider the modelling of an electoral competition. In an election there are a handful of key factors that are of relevance to a large number of electorates. These are the policy positions that would be taken by the candidates if they were elected, on taxation, public spending, abortion rights, gun control, climate policy, the Middle East, and so on. They may also include the personality or charisma of the candidates. Voters are unaware of which position a given candidate might take on any given issue, if elected, but have partial information about this. We let a random variable $X_{k}^{m}$ model the position of candidate $m$ on policy $k$.
Political messages will either be concerned with convincing electorates that their policy positions are advantageous to them, or saying whatever they think the electorates want to hear. Some messages will focus on convincing the electorates how the perceived policy positions of other candidates will be undesirable. In one way or another, there will be a lot of noise.
For each election factor $X^{m}_{k}$, we thus have a time series $\{\xi_{t}^{m,k}\}$ of noisy information, the aggregate of which for all $m$ and all $k$ determines the information available to the electorates, which they can use to arrive at their best estimates for these factors. Voters will have varied preferences — for instance, a position in support of stricter gun control may be perceived as good for one voter but bad for another. Voter preferences will be modelled as weighting factors that can be sampled from a distribution over voter preferences. Then for each voter we can assign scores for the candidates, with the rule that the candidate with the highest score on the election day will be selected by that voter. This is the idea of a structural model for an election (Brody & Meier 2002), which can be implemented in different practical situations.
Calculus for elections: a reduced-form approach
Although a structural model has advantages in micro-managing political messages during a major election campaign, qualitative features of such a model can be captured using a somewhat simplified reduced form model discussed now. Here we let a single random variable $X$ represent different candidates. Thus, if there are $n$ candidates then $X$ takes the values $x_{1},x_{2},\ldots,x_{n}$, labelling the candidates, with the a priori probabilities $p_{1},p_{2},\ldots,p_{n}$. These probabilities have the interpretation of representing the current poll statistics.
To understand the need of a dynamical model, consider an election involving two candidates, $A$ and $B$. Suppose that today’s poll shows that candidate $A$ has 52% support, and candidate $B$ has 48%; and the election is to take place in one year. Ignoring errors in poll statistics, and assuming a first past the post election system, what do these numbers tell us? What do they say about the likelihoods of the candidates winning the election?
What these numbers tell us is that if an election were to take place tomorrow, then the probability of candidate $A$ winning the election is about 100%. But they say very little about what happens in a future election. If the election were to take place in a year, then most people will think that the likelihood of candidate $A$ winning the election is close to 52%. But what if the election is to take place in six months, or in three months? To interpolate today’s statistics with future statistics, we need a dynamical model. In the present context, what generates the dynamics is the revelation of information, so let me take the simplest nontrivial model $\xi_{t}=\sigma Xt+\epsilon_{t}$ for the flow of information. Then we can generate a model for the dynamics of the poll statistics by the conditional probability $\pi_{it}={\mathbb{P}}(X=x_{i}|\xi_{\leq t})$, where $\xi_{\leq t}$ denotes information gathered up to time $t$. Using the Bayes formula we find, in this example,
$\pi_{it}=\frac{p_{i}\mbox{$\rm e$}^{\sigma x_{i}\xi_{t}-\frac{1}{2}\sigma^{2}x_{i}^{2}t}}{\sum_{i}p_{i}\mbox{$\rm e$}^{\sigma x_{i}\xi_{t}-\frac{1}{2}\sigma^{2}x_{i}^{2}t}}$
for the a posteriori support rates. This is how we can use a model for the flow of information to interpolate between today and the future. If the election is to take place at a future date $T$, then the vote share of candidate $j$ exceeding a threshold value $K$ on the election day, based on the knowledge today and on how information flows to the electorates from today till the election day, can be determined by working out the probability ${\mathbb{P}}(\pi_{jT}>K)$. As an example, in the case of a two-candidate race (Brody 2019), writing $p_{1}=p$ and $p_{2}=1-p$, a calculation shows that
${\mathbb{P}}\left(\pi_{1T}>K\right)=p\,N(d^{-})+(1-p)\,N(d^{+})\,$,
where
$d^{\pm}=\frac{\log\left(\frac{(1-p)(1-K)}{pK}\right)\pm\frac{1}{2}\sigma^{2}T}{\sigma\sqrt{T}}$
and $N(x)$ denotes the cumulative normal distribution function. It is a curious coincidence that the formula emerging from an information-based model for the projected probability ${\mathbb{P}}\left(\pi_{1T}>1/2\right)$, as of today, of a candidate winning a future election should coincide essentially with the famous Black-Scholes option pricing formula in asset pricing theory.
Figure 1: The probability of winning an election in one year, as a function of today’s support rate $p$, for $\sigma=0.2$ (purple) and $\sigma=1.2$ (red). When $p>0.5$, the realised probability of winning a two-candidate election in the future is always higher than $p$, and conversely for $p<0.5$.
There are lessons to be learned even from this simple model. If today’s polls were indicators of the likelihoods of candidates winning a future election, a common misconception, then we would have ${\mathbb{P}}\left(\pi_{1T}>1/2\right)=p$ for a two-candidate race. In reality the projected winning probability is always higher than today’s support rate $p$ if $p>1/2$; and always lower than $p$ if $p<1/2$. Exactly how much higher (or lower) depends on the information flow rate $\sigma$. For instance, suppose that one candidate has 52% support rate today, and that $\sigma\ll 1$. Then the realised probability of this candidate winning an election in a year is close to 1, because if that candidate has 52% support today, and if voters learn very little over the coming year because $\sigma\approx 0$, then those 52% of voters will still vote for this candidate, so the probability of winning is close to 1. This line of analysis shows that if you are leading in a two-candidate race then you want very little information to come out; whereas if you are lagging behind then you want as much information as possible to come out. In other words, why rock the boat when you are leading? This common phrase one hears from political commentators during an election cycle can be justified on account of a simple analysis in mathematical politics.
More generally, the rate $\sigma$ at which reliable information is revealed varies over time. In this case we lose the Markov property that the posterior probabilities are functions of the terminal value of the time series $\{\xi_{t}\}$. Nevertheless it is possible to work out the projected probabilities of the candidates winning a future election in closed form. By means of the method of functional derivative we can study optimal strategies for political messaging so as to maximise the winning probability (Brody 2019).
When there are three or more candidates in a race, such an analysis becomes nontrivial but nevertheless is tractable (Brody & Yuasa 2023). While in a two-candidate race the only control parameter is the signal-to-noise ratio $\sigma$, in an election with multiple candidates, their political spectrum gaps $x_{k}-k_{l}$ also become control parameters. What this means is that we can implement a sensitivity analysis to study how repositioning of the candidates or their political parties in the political spectrum will affect the winning probabilities. For example, suppose that there are three parties, positioning themselves at centre left, centre right, and right, and that the right-leaning party is ranked second in today’s poll. In this scenario, we find that if the right-leaning party moves further to the right, then their winning probability almost certainly drops, unless if $\sigma\ll 1$, that is, unless if the election is overshadowed by confusion and uncertainty. We thus encounter here another strength of mathematical analysis that explains the empirically observed fact that rightwing political parties tend to disseminate a lot of disinformation (Törnberg & Chueri 2025).
Tenacious Bayesians
Let me now discuss an important dynamical characteristic of the Bayesian updating leading to the transformation $p_{i}\to\pi_{it}$ from the prior to the posterior. For this purpose it will be useful to take the derivative of $\pi_{it}$, making use of the Ito stochastic calculus. Specifically, $\pi_{it}$ is a function of two variables $t$ and $\xi_{t}$, so its (stochastic) derivative $\mbox{$\rm d$}\pi_{it}$ can be worked out by expanding it as a Taylor series and retain terms of first order, bearing in mind that $(\mbox{$\rm d$}\xi_{t})^{2}=\mbox{$\rm d$}t$. If we then substitute $\mbox{$\rm d$}\xi_{t}=\mbox{$\rm d$}W_{t}+{\mathbb{E}}_{t}[\mbox{$\rm d$}\xi_{t}]$ into the result, we find that $\mbox{$\rm d$}\pi_{it}\propto\mbox{$\rm d$}W_{t}$. That is, increments of the posterior probability are proportional to those of the innovation process. Such a process is known as a ‘martingale’, representing a conserved quantity, albeit on average.
The stochastic differential equation satisfied by $\{\pi_{it}\}$ is known as the ‘Kushner equation’, which, in the present example is just an equation for a martingale, having no drift. To gain a deeper intuition of its dynamical characteristics, it will be helpful to bring in a geometric consideration. To proceed, consider the square root $\psi^{i}_{t}=\sqrt{\pi_{it}}$ of the posterior probability. Then squares of $\psi^{i}_{t}$ add up to unity, that is, $\{\psi^{i}_{t}\}$ determines a point on a unit sphere. The radius of this sphere, the total probability, however, is purely a matter of convention — we can say that the likelihood of an event is 0.6, or 6 out of 10, or 60%; they all convey the same notion. Statistically relevant information is thus encoded in the direction of the vector $\psi^{i}_{t}$, not the length. The space of directions in the vector space on which $\{\psi^{i}_{t}\}$ is defined is known as a ‘real projective space’, which is a compact Riemannian manifold $\mathfrak{M}$}. The dynamical equation for $\{\pi_{it}\}$ can then be lifted into a covariant differential equation (box below) on $\mathfrak{M}$}, given by
$\mbox{$\rm d$}\psi^{i}_{t}=-\textstyle{\frac{1}{16}}\nabla^{i}V_{t}\,\mbox{$\rm d$}t+\textstyle{\frac{1}{4}}\nabla^{i}X_{t}\,\mbox{$\rm d$}W_{t}.$
Here, $X_{t}(\psi)$ is the conditional mean and $V_{t}(\psi)$ is the conditional variance of the random variable $X$; both are functions on $\mathfrak{M}$}. The drift vector field of the state $\psi^{i}_{t}$, which determines the orientational direction, when there are three candidates, is plotted in figure 2.
Covariant Ito derivative (Hughston 1996, Brody & Hughston 2001)
An Ito process $\{x_{t}\}$ is defined by the equation $\mbox{$\rm d$}x_{t}=\mu_{t}\mbox{$\rm d$}t+\sigma_{t}\mbox{$\rm d$}W_{t}$, where $\{\mu_{t}\}$ and $\{\sigma_{t}\}$ are called the drift and the volatility of $\{x_{t}\}$. We can consider an Ito process $\{x_{t}\}$ taking values on a manifold $\mathfrak{M}$}, driven by an $m$-dimensional Wiener process $\{W_{t}^{i}\}_{i=1,\ldots,m}$. Let $\nabla_{a}$ be a torsion-free connection on $\mathfrak{M}$} such that for any vector field $\zeta^{a}$ its covariant derivative in local coordinates is $\delta_{\bf b}^{b}\delta_{a}^{\bf a}(\nabla_{b}\zeta^{a})=\frac{\partial\zeta^{\bf a}}{\partial x^{\bf b}}+\Gamma_{\bf bc}^{\bf a}\zeta^{\bf c},$ where $\delta_{a}^{\bf a}$ is the standard coordinate basis in a given coordinate patch. Suppose we have an Ito process taking values in $\mathfrak{M}$}. Let $x_{t}^{\bf a}$ denote the coordinates of the process in a particular patch. Then writing $h^{\bf ab}=\sigma_{i}^{\bf a}\sigma^{{\bf b}i}$ we define the drift process $\mu^{\bf a}$ by $\mu^{\bf a}\mbox{$\rm d$}t=\mbox{$\rm d$}x^{\bf a}+\mbox{$\textstyle\frac{1}{2}$}\Gamma_{\bf bc}^{\bf a}h^{\bf bc}\mbox{$\rm d$}t-\sigma_{i}^{\bf a}\mbox{$\rm d$}W_{t}^{i}$. Alternatively, we can write the covariant Ito differential as $\mbox{$\rm d$}x^{a}=\delta_{\bf a}^{a}(\mbox{$\rm d$}x^{\bf a}+\mbox{$\textstyle\frac{1}{2}$}\Gamma_{\bf bc}^{\bf a}h^{\bf bc}\mbox{$\rm d$}t)$, where $\delta^{a}_{\bf a}$ is the dual coordinate basis. Then we have $\mbox{$\rm d$}x^{a}=\mu^{a}\mbox{$\rm d$}t+\sigma^{a}_{i}\mbox{$\rm d$}W^{i}_{t}$ for the covariant Ito differential on $\mathfrak{M}$} associated with the given connection.
Figure 2: Negative gradient flow of the variance. The three corners of the octant, where the flow attempts to take you to, correspond to states with zero uncertainty.
The important point is that while $\pi_{it}$ has no drift, the induced process on $\mathfrak{M}$} has a drift that is given by the negative gradient of the variance of $X$. In other words, there is a tendency to flow towards the directions with zero uncertainty about the choice in question.
What this means is that in assessing noisy information, people are drawn into states with lower uncertainties. In an extreme situation whereby a person is almost certain that one of the alternatives is the right choice, this state of zero uncertainty acts as a trap. That is, even if the person obtains noisy information that suggests that this is the wrong choice, a rational person who follows Bayesian updates will hardly change their view on the matter for a very long time (figure 3). I called this effect a ‘tenacious Bayesian’ behaviour (Brody 2022).
Today’s politics plays out against a backdrop of uncertainties that include wars in Ukraine and Gaza; the cost of living crisis; energy, food and water insecurity; migration; and so on. Above all, the impact of the climate crisis. The electorates are more than ever longing for certainties. This is why they are drawn into candidates who can offer a sense of certainty, irrespective of whether it is true or false. The rise of ‘populist’ parties around the world, who offer simple solutions to complex societal challenges, when there are no simple solutions available, is a manifestation of this effect. Failing to realise the importance of this effect — an effect that becomes clearly visible when applying stochastic differential geometry — in political messaging, a candidate is unlikely to succeed in an election.
Modelling the impact of ‘fake news’
Figure 3: If $p_{1}=0.01$ and $p_{2}=0.99$, then even if the noisy information is provided by $\xi_{t}=\sigma x_{1}t+\epsilon_{t}$, indicating that $X=x_{1}$ is the correct alternative so that $\pi_{1t}$ should approach 11, a Bayesian thinker will maintain the position $\pi_{1t}\approx 0$ for a long time. Four sample paths are shown here for $\sigma=1$.
Because we have a formalism to model how the flow of information affects people’s behaviour, we can include a disinformation term to study its impact. To this end, note that a genuine noise (rumours, speculations, and so on) carries no unknown bias so that we can assume ${\mathbb{E}}[\epsilon_{t}]=0$, because any known noise bias will be discarded. Hence if there is an unknown bias in the noise, this will skew the assessments of the receiver of the information. With this in mind, a simple model for disinformation can take the form
$\xi_{t}=\sigma Xt+(\epsilon_{t}+f_{t})\,$,
where ${\mathbb{E}}[f_{t}]\neq 0$ and where the existence of the term $f_{t}$ is unknown to the receiver of the message.
With the help of the model we can conduct a quantitative statistical analysis of the impacts of disinformation on an election. For example, we can determine the averaged percentage change in the election outcome if disinformation of a given magnitude and frequency were disseminated. One result that might come as a surprise is that if the receiver is aware of the statistical distribution of the term $f_{t}$, then this is sufficient to almost entirely eradicate its impact (Brody & Meier 2022). Fact checkers are helpful in determining whether any given information is true or false. But a more useful information is the detailed statistics of their findings, which is not commonly available. Mathematical analysis shows that such information is helpful to counter the impacts of disinformation in democratic processes, but currently, policy makers seem unaware of the fact that such mathematical analysis can be applied to develop countermeasures to tackle the issue.
There is another situation in which an effective countermeasure can be implemented when right conditions are met. This concerns how information are combined. If there are two information sources, $\xi_{t}^{1}=\sigma_{1}Xt+\epsilon_{t}^{1}$ and $\xi_{t}^{2}=\sigma_{2}Xt+\epsilon^{2}_{t}$, on the same issue $X$, then the combined effect is equivalent to the consumption of a single information source
$\xi_{t}=\frac{\sigma_{1}-\rho\sigma_{2}}{\sigma(1-\rho^{2})}\,\xi_{t}^{1}+\frac{\sigma_{2}-\rho\sigma_{1}}{\sigma(1-\rho^{2})}\,\xi_{t}^{2}\,$,
where $\rho$ is the correlation of the two Brownian noises $\epsilon^{1}_{t}$ and $\epsilon^{2}_{t}$, and $\sigma$ is a certain positive function of the parameters $\sigma_{1}$, $\sigma_{2}$, and $\rho$. We can think of $\xi^{1}_{t}$ and $\xi^{2}_{t}$ representing, for instance, information emerging from two competing political parties. Suppose that $\rho>0$ and that $\xi_{t}^{2}$, say, contains disinformation. In this case, by letting $\sigma_{1}>\sigma_{2}/\rho$, it is possible to reverse the impact of disinformation, that is, fake news will turn against its spreader (Brody & Yuasa 2025). The lesson for the first political party therefore is to deliver clear and transparent messages to enhance $\sigma_{1}$, rather than to pay too much attention to counter the disinformation.
A pathological liar?
What if you have a prime minister, or a president, who is a pathological liar? I shall not get into clinical definitions of a pathological liar here, but instead interpret it to simply mean a person who knows the fact but is determined not to reveal it for one reason or another. With this understanding, can mathematics uncover statistical signs of the behaviour of such a liar? The information-based approach is sufficiently versatile to give us an affirmative answer.
Figure 4: There are six alternatives labelled by $X=1,2,3,4,5,6$. If $X=3$ but $p_{3}=0$, then the mean $X_{t}$ tends to hop between the closest alternatives $X=2$ and $X=4$, when there is a lot of ($\sigma=5$) reliable information coming out saying that $X=3$ (top panel). In contrast if $p_{3}=0.01$ but $X=3$, a genuine misunderstanding, then $X=2$ and $X=4$ remain plausible alternatives for a while, but the view will gradually converge to the true alternative $X=3$ (bottom panel). Six sample paths are shown in each case.
Recall that the posterior probabilities $\pi_{it}=\pi_{i}(t,\xi_{t})$ are functions of two variables, time $t$ and information $\xi_{t}=\sigma Xt+\epsilon_{t}$. One of the advantages of the information-based approach is that numerical simulations are very easy to implement because there is no need to solve stochastic differential equations. Specifically, we sample a value of $X$ and a Brownian path $\epsilon_{t}$, substitute them in $\xi_{t}$ and plug that into $\pi_{i}(t,\xi_{t})$ to generate a simulation outcome, for example, of the expected mean $X_{t}=\sum_{i}x_{i}\pi_{it}$. The sampling of $X$ is done according to the prior probabilities $\{p_{i}\}$.
Now suppose that $p_{k}=0$ for one of the alternatives. It is nonetheless possible to consider the sample $X=x_{k}$; an event that occurs with zero probability, and study the behaviour of the conditional mean $X_{t}$. When there is a lot of reliable information coming out saying that $X=x_{k}$ is the correct alternative, the result is quite revealing. The process $X_{t}$ rather abruptly converges to one of the ‘false’ alternatives $x_{l}$, $l\neq k$, and stays there for a while. But noisy information being revealed suggests that $X=x_{k}$, so eventually $X_{t}$ is kicked off from the level $x_{l}$ and quickly converges into another false alternative $x_{m}$, and repeat this procedure by hopping from one false alternative to another, typically between the closest false alternatives $x_{k+1}$ and $x_{k-1}$ to the true alternative $x_{k}$ (figure 4).
In other words, a pathological liar of the kind defined above would quite confidently pick one of the false alternatives and maintains the narrative, until enough information is revealed to make that position uncomfortable. At this point a different false alternative is quickly and confidently chosen, as if there is no inconsistency with the previous position, until when he is forced to change his position again.
The behaviour of the process $\{X_{t}\}$, which is something reminiscent of a piecewise step function, when $p_{k}=0$ but $X=x_{k}$, is statistically distinguishable from the behaviour of $\{X_{t}\}$ when $p_{k}\ll 1$ but is nonzero, no matter how small $p_{k}$ may be. The latter represents a situation in which a person is genuinely unaware that $X=x_{k}$ is the correct choice. So from the behaviour it is possible, at least with some statistical confidence, to determine whether someone is refusing to admit the truth (by blocking out the true alternative $X=x_{k}$) or is genuinely unaware of the truth, in an environment where noisy information about the truth is revealed.
Transmission, feedback, and control
Thus far our attention has been focused on how noisy information is processed and analysed, which we used to deduce the resulting action. Given the fact that this analysis generates an uncertainty-minimising flow, it is natural to ask whether an individual can transmit information back into the system in the form of a feedback, with the view to further reduce uncertainties. That is, we can think of the possibility of feedback control. To this end consider an individual (the system), the environment (community, society, or a section of the society), and the communication channel (Internet, Twitter / X, Instagram, TikTok, and so on).
It is a well-known fact that channel capacity cannot be increased by feedback. Nevertheless, it turns out that noise-free feedback from the system to the environment can enhance the information gained from the environment. Consider the following, somewhat idealised setup of Davis (1978). In this setup, the information entering the system from the environment is obscured by noise, but the environment can detect the signal transmitted from the system without noise. The system (individual) is interested in the amount of information contained in the message coming from the environment about the value of the signal. This quantifier is known in communication theory as mutual information (Gelfand & Yaglom 1957), which is determined by the likelihood ratio for detecting the signal. The question therefore is what feedback should the individual transmit back into the community so as to maximise the mutual information. In the case of a Gaussian signal, the optimal transmission turns out to be the increments $\mbox{$\rm d$}W_{t}$ of the innovation process itself (Davis 1978). That is, only new information is transmitted from the system to the environment to maximally reduce the uncertainty of the signal.
There is a profound implication of this result, albeit in a slightly idealised setup, in today’s digital society. From individual’s perspective, transmission of information is time and energy consuming. Hence in practice unless the magnitude $|\mbox{$\rm d$}W_{t}|$ of the new information is above a threshold, it is unlikely that this is fed back into the community. Recall that the new information $\mbox{$\rm d$}W_{t}$ is the discrepancy $\mbox{$\rm d$}\xi_{t}-{\mathbb{E}}_{t}[\mbox{$\rm d$}\xi_{t}]$ between the predicted and realised increments of the information. Now suppose that the information $\mbox{$\rm d$}\xi_{t}$ over the time interval $t,t+\mbox{$\rm d$}t]$ is secretly contaminated with disinformation $\Delta f$. Then this will shift the realised increment $\mbox{$\rm d$}\xi_{t}$ by $\Delta f$ but not the predicted increment ${\mathbb{E}}_{t}[\mbox{$\rm d$}\xi_{t}]$, because the receiver is unaware of the existence of disinformation. It follows that for sufficiently large $\Delta f$, the likelihood of the distorted innovation $|\mbox{$\rm d$}W_{t}+\Delta f|$ exceeding the threshold value is very high.
It is known empirically that the likelihood of a person disseminating disinformation is significantly higher than that of the person disseminating true information (Vosoughi et al. 2018). The observation here demonstrates the intuitively acknowledged fact that the higher the novelty of information, the more likely it is that the information is transmitted. Importantly, though, this is the consequence of a highly rational behaviour, following the rule of optimal feedback control. The consequence therefore is to further enhance noise in the system, and this enhancement is amplified in a society that is becoming more reliant on the use of artificial intelligence systems.
Political polarisation
The tenacious Bayesian behaviour shows that if a person is rational in the sense of processing information in accordance with the Bayes formula, then in the event of that person believing strongly in a false alternative, the scope of changing the mind of that person by supplying factual information is limited. It is then reasonable to ask whether an alternative approach might exist to deal with such circumstances whereby misguided people holding strong but undesirable political views.
It is often pointed out that politics is becoming more and more polarised, but the same is applicable to society as a whole. To understand this phenomenon, it will be useful to cite Wiener’s argument (Wiener 1950) that the effects resulting from the second law of thermodynamics on society cannot be overstated. The second law says that over time, order will be replaced by confusion and uncertainty – a phenomenon that is inevitable in a society with rapidly-developing communication technology.
A rational individual, on the other hand, seeks order and certainty. Indeed, one can show that the Bayesian updating follows from the requirement of maximal reduction in the expected uncertainty (Brody & Trewavas 2023). That is, when we acquire noisy information $\xi$ about a quantity of interest $X$, there are in general uncountably many transformations from the prior view $p_{i}$ to the posterior view $\pi_{i}$ that are all consistent with the acquired information. But if we demand that the transformation should minimise, on average, the uncertainty of $X$, then the result agrees with the expression obtained by use of the Bayes formula for the conditional probability. Biologically this makes a lot of sense. If $X$ represents the unknown condition of the environment, for instance, then a protocol that maximally reduces uncertainties of $X$ is highly advantageous for survival because a biological system must adapt to the condition of its environment. Thus we have a conflict: The entropy that measures the degree of uncertainty of the society is rapidly increasing, but individually we try to reduce entropy. A corollary is that to keep the level of entropy low in such an environment, a person must limit the latitude of information source, resulting in the creation of echo chambers. The phenomenon of social polarisation therefore is the result in part of the competition between the second law and the individual desire to counter the second law. It should be evident that recent advances in artificial intelligence can only enhance this conflict.
The situation is further exacerbated by the fact that unlike in mathematics or in physical science where the validity of a proposition can be tested, either by a mathematical proof or by repeatable experiments, in politics or in social science more generally, this is often not possible. The long-term consequence of an economic policy on society, for instance, cannot be determined with certainty. One might have an opinion on the matter, but it remains an opinion, no matter how informed that opinion might be. Hence the boundary of fact and fallacy is a fuzzy one in politics, making it harder for even a well-informed opinion to be heard by political opponents.
Incompatible propositions
Given that the effect of the second law cannot be eradicated, the issue of political polarisation is a challenging one to address. Nevertheless, when there is an unambiguous fact available, for instance in the case of an event that had actually taken place, it is legitimate to ask whether a misguided individual strongly believing in a false alternative can be influenced.
To explore this possibility, it is worth remarking that the Bayes formula, which tells us how a prior view on a topic is updated in accordance with the acquired information, is applicable only if the prior and the posterior probabilities are absolutely continuous with respect to each other, in the sense that they share a common null set. When the continuity condition is lost, or more generally when a sequence of information variables does not admit a joint probability, the Bayes formula is not applicable. However, even in those circumstances, an uncertainty-minimising transformation does exist, and is given by what is known as the Lüders transformation. The idea can most easily be explained by relying on geometric intuitions and consider the square roots $\sqrt{p_{i}}$ and $\sqrt{\pi_{i}}$ of the prior and posterior probabilities.
The vector space in which square-root probabilities are defined, when endowed with a Euclidean inner product, is known as a Hilbert space. This is the Hilbert space ${\cal H}_{\cal S}$ that models our view of the ‘signal’ $X$. There is likewise a Hilbert space ${\cal H}_{\cal N}$ that models our view of the ‘noise’ $\epsilon$. The tensor-product space ${\cal H}_{\cal S}\otimes{\cal H}_{\cal N}$ then models the space of all possible square-root joint probabilities of $X$ and $\epsilon$, where random variables on Hilbert space are modelled by matrices. When a noisy information is acquired, that is, when the value of $\xi=\sigma X+\epsilon$ is detected, this singles out a linear subspace ${\cal H}_{\xi}$ of ${\cal H}_{\cal S}\otimes{\cal H}_{\cal N}$ corresponding to all possible combinations of the random variable pair $(X,\epsilon)$ such that the value of $\sigma X+\epsilon$ agrees with the detected value of $\xi$. The situation is best understood by considering the case in which $X$ is discrete and $\epsilon$ is continuous. Then given the value of $\xi=y$, we know either $X=x_{1}$ and $\epsilon=y-\sigma x_{1}$, or $X=x_{2}$ and $\epsilon=y-\sigma x_{2}$, or $X=x_{3}$ and $\epsilon=y-\sigma x_{3}$, and so on — all these possibilities are consistent with the data $\xi=y$. The idea of the Lüders rule is that the square-root of the a priori joint probability is projected down onto ${\cal H}_{\xi}$ to form the square-root of the a posteriori joint probability. When the result obtained is squared, and suitably normalised, we recover the Bayes formula. The geometric formulation thus uncovers the fact that not only the transformation from the prior to the posterior probabilities maximally reduces uncertainties but also it generates the minimum change in the probability that is consistent with the acquired information (Brody & Trewavas 2023). This, again, is biologically advantageous because the smallest change in the view of the external world that is consistent with detected data leads to the smallest waist in energy consumption.
Now suppose that having acquired the information $\xi=\sigma X+\epsilon$ one subsequently acquires another information, which is modelled by $\xi^{\prime}=\sigma Y+\epsilon^{\prime}$, where the two random variables $X$ and $Y$ do not admit a joint probability. In this case, the Bayes formula is not applicable, whereas the Lüders projection rule is always applicable. When the two random variables $X$ and $Y$ do not admit a joint probability, we say that the two propositions modelled by these random variables are not compatible. Practically, what this means is that the matrix representations of $X$ and $Y$ on Hilbert space are not commutative. Thus we enter into the domain of noncommutative communication theory. Such a seemingly abstract concept turns out to be highly relevant in political messaging.
In cognitive psychology, for long it was taken for granted, quite understandably, that all propositions are compatible. One consequence of this assumption is that when people’s behaviour deviates away from what is predicted by the rules of probability, for example a violation of the Bayesian updating, such a behaviour can be classified as irrational. A classic example is the notorious conjunction fallacy (Tversky & Kahneman 1983) , whereby “$A$ and BB” is thought more likely than“AA”. According to standard logic and probability theory this is never true, but in experiment after experiment it can be human-true. But standard logic and probability theory is applicable only so far as when $A$ and $B$ admit a joint probability law.
It is legitimate therefore to ask: What if $A$ and $B$ were not compatible? This is in effect the question addressed by Jerome Busemeyer and others in cognitive science (see Busemeyer & Bruza 2024, and references cited therein). The answer they have found was remarkable. Much of the seemingly irrational human behaviours reported in the cognitive psychology literature, for which there are no straightforward explanations using only compatible propositions, could be explained consistently using models accommodating incompatible propositions. Statistical analysis of random events involving incompatible propositions, however, is what quantum theory is about. Indeed, there are two fundamental distinctions between Kolmogorovean classical probability theory and what is known as quantum probability theory. The first is that the square-root probabilities $\psi_{i}$, called probability amplitudes in quantum mechanics, can take complex values so that the probabilities are recovered using modulus squares: $p_{i}=|\psi_{i}|^{2}$. (Imaginary number plays two roles in quantum mechanics, namely, to rotate plane waves by the right angle, and to set the direction of time. It is known that quantum theory cannot be formulated without the use of complex numbers, whereas it is not known whether complex numbers are required for a consistent characterisation of cognitive behaviours.) The second is that random variables, which are modelled on Hilbert space as matrices, are in general not compatible. With the realisation that quantum probability models can more accurately and succinctly model cognitive behaviours, a new field emerged within cognitive science called ‘quantum cognition’, which is now an active field of research.
It may be added parenthetically that quantum cognition is not concerned with quantum effects that may or may not be taking place at the neurophysiological level. Rather, it concerns the application of quantum probability assignment rules to model cognitive human behaviour. If a pattern of human behaviour appears to violate rules of classical probability but respects quantum rules, then one can no longer classify this as being irrational, because when there are incompatible propositions, classical probability is simply not applicable to say what the rules are. The situation is entirely analogous to the discussion concerning the famous Bell inequality in quantum mechanics (Isham 1995). The bound on a certain correlation in the Bell inequality is obtained by assuming the existence of a joint probability distribution, so when there is no such distribution, classical probability is simply not applicable. It is then illogical to argue that a violation of the Bell inequality is a violation of classical laws of physics, because classical physics provides no bound. Where it violates classical physics is the existence of incompatible propositions itself, and not the inequality. In a similar vein, if a seemingly irrational behaviour is consistent with the rules of quantum probability, then this should be viewed as entirely rational, because quantum rules provide a high level of efficiency that should be preferred biologically (Brody & Trewavas 2023).
The relevance of what I call here noncommutative communication theory in politics has long been appreciated by seasoned politicians and political advisors, purely on the empirical or intuitive grounds, as an art of messaging. Perhaps a simple anecdotal example suffices to highlight this. Take the 2016 ‘Brexit’ referendum in the United Kingdom on membership of the European Union. Some ‘leave’ campaigners put to the public the question, before the referendum, whether they want more funding for the National Health Service, with the implication, true or false, that leaving the European Union leads to more funding for the health service. If the membership question and the health service funding question were compatible, then classical probability theory shows that asking the funding question will have no implication in the referendum outcome. Those leave campaigners believed otherwise, and the result was ‘out’.
There are innumerable such examples whereby political campaigners instinctively know how to shift the frame of reference of the public opinion, but the science underpinning this effect is only emerging in the form of quantum cognition and noncommutative communication theory, over the past two decades. Indeed, even in the hypothetical scenario whereby a person believes 100% in an alternative, there is a way of changing the view of that person – an impossibility in classical probability (Basieva et al. 2017). The trick is not to convince the person otherwise, which is futile because a counterargument to that alternative is compatible with the argument itself. Counterintuitive this may be, a counterargument will have no effect on the view of that person. Instead, by asking a single incompatible question, and importantly a single question is sufficient here, it is possible to change the belief of the person on the original issue to a number strictly less than 100%.
Discussion
As a discipline, mathematical politics in the sense of an art of messaging is very much at the stage of infancy. While much insights have been gained with perhaps the simplest nontrivial models, there are scopes for further developments. Indeed, mathematical theory underpinning signal detection, sometimes called ‘stochastic filtering theory’, has developed considerably over the past 60 years. It seems inevitable that some of the more powerful machineries can be applied to gain a wide range of new insights in political messaging and beyond.
There are new mathematical challenges that emerge as well. For example, in traditional communication theory, there are no ambiguities as for the interpretation of the signal (the random variable $X$ in the current example). In politics, however, there is a huge ambiguity. Consider a newspaper article on an issue. The view of a person on this issue may be modelled by $X$ (taking the values $\{x_{i}\}$ with the probabilities $\{p_{i}\}$) so that the information contained in the article can be modelled, say, by $\xi=\sigma X+\epsilon$; but only for that person. Another reader of the same article will have a different view on the issue, represented by $Y$ (also taking the values $\{x_{i}\}$ but with the probabilities $\{q_{i}\}$), so the information model for that person takes the form, say, $\xi^{\prime}=\sigma Y+\epsilon$, even though we are speaking about an identical news article. Thus, mutual information of a message, that is, the amount of information contained in the message about the value of the signal, is receiver dependent. In a time series version of the problem, statistical analysis requires the specification of information in the form of conditioning, but when there are ambiguities regarding which, or whose information to consider, it is not clear how to conduct such analysis, and this is a challenge already in pure mathematics.
Beyond politics, the programme introduced here can have profound implications in other areas. One is in cognitive psychology. For instance, the tenacious Bayesian phenomenon discussed above can be used to show that some of the common conclusions in the psychology literature on the seemingly irrational behaviour of people are not always correct (Brody 2022). Mathematical analysis, when applied properly, offers so much potential to model and explain cognitive behaviours, both qualitatively and quantitatively. When this is combined with developments in quantum cognition and noncommutative communication theory, the scope of a better understanding of human behaviour, in the cognitive context, broadens further. At the same time, one should not underestimate the risks involved, for, a better understanding of our cognitive response to signals can potentially be used to control people, both in positive and negative ways.
Another application outside politics is in advertisement and marketing. Because the mathematical theory of politics is concerned with the understanding of the impacts of political messaging, the findings can be extended, mutatis mutandis, to marketing. This is because the questions being addressed are very similar in the two fields — how to maximise the impact of a message, either in the form of vote share or of the sales figure. In both areas, more traditional regression analysis is commonly employed, but the mathematical theory of information and communication can offer considerably more towards generating future statistics.
It is of interest also to explore behaviours of artificial intelligence from the viewpoint of the present discussion. The mathematics of messaging is concerned with understanding how a signal is processed by an intelligent system. Presumably today’s artificial intelligence would process signals differently from what the theory predicts, but if so, how different are they? It is hoped that a better understanding of the mathematics of messaging will lead to an alternative and better design of an artificial intelligence architecture that is not only transparent but also is significantly more energy efficient — an endeavour that is not too dissimilar to the attempts based on the framework of active inference (Pezzulo et al. 2024).
The idea of modelling information surrounding us as a probabilistic time series may at first seem abstract and intangible. As Edgar Allan Poe wrote, “Such sentiments are seldom thoroughly stifled unless by reference to the doctrine of chance, or, as it is technically termed, the Calculus of Probabilities. Now this Calculus is, in its essence, purely mathematical; and thus we have the anomaly of the most rigidly exact in science applied to the shadows and spirituality of the most intangible in speculation.” The information-based approach to mathematical politics shows that the matter is not quite so intangible as one might have thought at first.